OpenStudy (anonymous):

Simplify : 11/12 ÷( 2/3 + 7/9)

OpenStudy (zephyr141):

find the common denominator of 2/3 and 7/9 which is 9 so yay it's easy. multiply 2/3 by three and that leaves us with 6/9. now you can add the two terms and you would get \[\frac{ 6+7 }{ 9 }=\frac{ 13 }{ 9 }\]

OpenStudy (zephyr141):

now to finish this you can divide \[\frac{ \frac{ 11 }{ 12 } }{ \frac{ 13 }{ 9 } }\] or you can multiply instead by using the reciprocal of 13/9 which is just 9/13 \[\frac{ 11 }{ 12 }*\frac{ 9 }{ 13 }=?\] that would be simplifying it entirely.

OpenStudy (anonymous):

1089/156

OpenStudy (zephyr141):

\[\frac{ 11*9 }{ 12*13 }\] sorry for the delay. the answer is \[\frac{ 99 }{ 156 }\]